A tank contains a mixture of 48.2 g of oxygen gas and 70.0 g of carbon dioxide gas at 23 °C. The total pressure in the tank iWhat mass of neon gas is required to fill a 8.00 L container to a pressure of 1.42 atm at 18 °C? Mass = g NeHow many grams of mercury(II) oxide are needed to produce 4.04 L of oxygen gas according to the following reaction at 25 °C a
Answer #1

1)

Molar mass of O2 = 32 g/mol

Molar mass of CO2,

MM = 1*MM(C) + 2*MM(O)

= 1*12.01 + 2*16.0

= 44.01 g/mol

n(O2) = mass of O2/molar mass of O2

= 48.2/32

= 1.5063

n(CO2) = mass of CO2/molar mass of CO2

= 70.0/44.01

= 1.5905

n(O2),n1 = 1.5063 mol

n(CO2),n2 = 1.5905 mol

Total number of mol = n1+n2

= 1.5063 + 1.5905

= 3.0968 mol

Partial pressure of each components are

p(O2),p1 = (n1*Ptotal)/total mol

= (1.5063 * 8.4)/3.0968

= 4.0857 atm

p(CO2),p2 = (n2*Ptotal)/total mol

= (1.5905 * 8.4)/3.0968

= 4.3143 atm

Answer:

p(O2) = 4.09 atm

p(CO2) = 4.31 atm

Only 1 question at a time please